OpenStudy (anonymous):

Please please please help!!! Explain how to find the average rate of change for f(x) from x=4 to x=7 I have equation x^2-4x+12

OpenStudy (imstuck):

You need to plug 4 in for x and find the y; then you need to plug 7 in for x and find THAT y. Then do the slope formula, y2 - y1/x2 - x1 to find the rate of change.

OpenStudy (anonymous):

When I plug them in I got 12 and 33. Is that correct so far?

OpenStudy (anonymous):

\[\text{average rate of change over }(a,b)=\frac{f(b)-f(a)}{b-a}\] where \(a=4\) and \(b=7\). You have \[\begin{cases}f(7)=7^2-4(7)+12=33\\ f(4)=4^2-4(4)+12=12\end{cases}\] So from here, the avg rate of change would be \[\frac{33-12}{7-4}=\cdots\]

OpenStudy (anonymous):

You would get 21/3 Which would be 7. so your Average rate of change is 7 c:

OpenStudy (anonymous):

Ohhhh ok that seems simple! Thank you guys so so much! Btw, sorry I lost internet for a while.